TR = 42Q -0.5Q2 TC=Q3-8Q2+57Q+50 Find the profit maximizing Output level.
Answer #1

An TR = 429 -0.59² Revenue cost Tc = q3 - 892 +579 + 50 Profit maximization condition : MR=MC where MR=marginal MC = Marginal